A particle is described by a position Q (2.5, π/3, 6) in a cylindrical coordinate, determine the position of the particular Cartesian coordinate


1
Expert's answer
2022-01-25T06:50:33-0500

To determine the position of the particle in Cartesian coordinates, apply the following transformations:


"x=r\cos\phi=2.5·\cos\dfrac\pi3=1.25,\\\space\\
y=r\sin\phi=2.5·\sin\dfrac\pi3=2.17,\\\space\\
z=z=6."